# Auditing Algorithmic Bias in Transformer-Based Trading

### v1

@@ 1 Introduction
The recent advancements of advanced artificial intelligence is instigating a paradigm shift in quantitative finance. The auto-regressive Transformer models Vaswani et al. [2017], which have powered recent breakthroughs in natural language processing and computer vision, are now being rapidly adopted for algorithmic stock trading Coelho e Silva et al. [2024]. Architectures such as the "Quantformer" Zhang et al. [2024] leverage sophisticated self-attention mechanisms to capture complex, long-range dependencies within noisy financial time-series, demonstrating empirical superiority over predecessor models like Long Short-Term Memory (LSTM) networks. These models achieve higher predictive accuracy and generate superior cumulative returns, establishing a new state-of-the-art in financial forecasting Guo et al. [2025]. However, their complexity gives rise to a critical challenge: the models operate as opaque "black boxes," where the intricate path from data input to trading decision is largely an enigma, even to their developers Umeaduma [2025]. This lack of transparency poses significant risks for accountability, fairness, and regulatory compliance, creating a need for rigorous auditing methodologies.
@@ 2.2 Partial Information Decomposition
PID Kolchinsky [2022] is a theoretical framework designed to analyze how a set of input variables provides information about an output variable. Its goal is to move beyond simple mutual information and dissect how the information is structured among the sources. The central idea is to break down the total information into distinct, non-overlapping components, each describing a different mode of interaction:
@@ 3.1 Training Transformer
The network has an embedding size of $D=256$, $H=8$ attention-heads per attention layer, and four attention layers. Furthermore, we use ROPE Pochet et al. [2023] for positional encoding, and take advantage of add-and-norm residual connection. We set the context length to $N=64$, meaning the model uses the $pr$ of the target and support stocks from the previous $64$ timesteps to predict the target’s $pr$ for the next timestep. We use mean Cross-Entropy as our loss function, and apply a dropout rate of $0.1$ to prevent overfitting.
@@ 4.1 Reliance Independent of IV
Relying on a support stock with high IV is risky since its price is more likely to experience large, sudden swings Addy et al. [2024]. These movements can influence the model, causing its prediction to mirror the support stock’s trajectory. If this movement is specific to that stock and not reflective of the broader market, it can lead the model to predict an incorrect $pr$ direction. For example, the model relies most heavily on AMD (as shown by its low excluded information) despite it having the highest IV. If company-specific news were to cause a sharp drop in AMD’s price, this high reliance would skew the model’s prediction for NVDA downward, even if the news had no bearing on NVDA’s actual value.

### v2

@@ 1 Introduction
The recent advancements of advanced artificial intelligence is instigating a paradigm shift in quantitative finance. The auto-regressive Transformer models vaswani2017attention, which have powered recent breakthroughs in natural language processing and computer vision, are now being rapidly adopted for algorithmic stock trading coelho2024transformers. Architectures such as the "Quantformer" zhang2024quantformer leverage sophisticated self-attention mechanisms to capture complex, long-range dependencies within noisy financial time-series, demonstrating empirical superiority over predecessor models like Long Short-Term Memory (LSTM) networks. These models achieve higher predictive accuracy and generate superior cumulative returns, establishing a new state-of-the-art in financial forecasting guo2025trading. However, their complexity gives rise to a critical challenge: the models operate as opaque "black boxes," where the intricate path from data input to trading decision is largely an enigma, even to their developers umeaduma2025explainable. This lack of transparency poses significant risks for accountability, fairness, and regulatory compliance, creating a need for rigorous auditing methodologies.
@@ 2.2 Partial Information Decomposition
PID kolchinsky2022novel is a theoretical framework designed to analyze how a set of input variables provides information about an output variable. Its goal is to move beyond simple mutual information and dissect how the information is structured among the sources. The central idea is to break down the total information into distinct, non-overlapping components, each describing a different mode of interaction:
@@ 3.1 Training Transformer
The network has an embedding size of $D=256$, $H=8$ attention-heads per attention layer, and four attention layers. Furthermore, we use ROPE pochet2023roformer for positional encoding, and take advantage of add-and-norm residual connection. We set the context length to $N=64$, meaning the model uses the $pr$ of the target and support stocks from the previous $64$ timesteps to predict the target’s $pr$ for the next timestep. We use mean Cross-Entropy as our loss function, and apply a dropout rate of $0.1$ to prevent overfitting.
@@ 4.1 Reliance Independent of IV
Relying on a support stock with high IV is risky since its price is more likely to experience large, sudden swings addy2024algorithmic. These movements can influence the model, causing its prediction to mirror the support stock’s trajectory. If this movement is specific to that stock and not reflective of the broader market, it can lead the model to predict an incorrect $pr$ direction. For example, the model relies most heavily on AMD (as shown by its low excluded information) despite it having the highest IV. If company-specific news were to cause a sharp drop in AMD’s price, this high reliance would skew the model’s prediction for NVDA downward, even if the news had no bearing on NVDA’s actual value.